OpenStudy (anonymous):

Give a unique example of each: An absolute value equation with two solutions An absolute value equation with one solution An absolute value equation with no solutions

OpenStudy (aripotta):

\[\left| x + 2 \right| = 7\]\[-\left| x \right| = 9\]\[\left| x - 4 \right| = -14\]

OpenStudy (aripotta):

ugh wait. that second one is wrong i think

OpenStudy (aripotta):

\[\left| x \right| = 0\] i guess that would work